OpenStudy (mtalhahassan2):

Perform the following operations to the correct number of significant figures. 257 N X 0.42 m =

OpenStudy (mtalhahassan2):

@Shalante

OpenStudy (jhannybean):

When multiplying 2 numbers take the least number of significant figures. how many sig figs does 257 have? how many dig figs does 0.42 have?

OpenStudy (mtalhahassan2):

257 have 3 and 0.42 have 2

OpenStudy (mtalhahassan2):

so we are goona use two s.f right??

OpenStudy (mtalhahassan2):

which be 107.94

OpenStudy (jhannybean):

Nope. 107.94 = 5 sig figs. All the numbers count, even the 0 in the middle of the 1 and 7.

OpenStudy (mtalhahassan2):

but know here we have 5 s.f right

OpenStudy (mtalhahassan2):

then it be 1.1X10^2

OpenStudy (jhannybean):

You are correct

OpenStudy (mtalhahassan2):

so this be the answer

OpenStudy (jhannybean):

mmhmm.\(\checkmark\)

OpenStudy (mtalhahassan2):

but one of my friend said that it be 110 Newton meters

OpenStudy (mtalhahassan2):

how is it comes thou??

OpenStudy (mtalhahassan2):

is they convert newton to meters maybe

OpenStudy (jhannybean):

110 = 3 sig figs which is incorrect.

OpenStudy (mtalhahassan2):

no its two because the last zero doesn't count

OpenStudy (jhannybean):

Oh my bad. Yes, you are right, 110 has 2 sig figs, but 110. has 3

OpenStudy (mtalhahassan2):

wait but how he get 110

OpenStudy (mtalhahassan2):

@Shalante

OpenStudy (jhannybean):

I `think` he got 110 by rounding 107.94 to 2 decimal places... so rounding 0 to 1 without touching the decimal. Therefore it would become 110. I believe it is wrong because that is excluding all the other numbers as part of the answer, whereas \(\sf 1.1 ~\times~ 10^2\) does not disregard the other numbers, it just states that \(\sf 1.1~\times~10^2\)is a more accurate representation of the answer to 2 significant figures.

OpenStudy (mtalhahassan2):

1.82 g/mL X 0.613 cm = 1.6X10^1

OpenStudy (mtalhahassan2):

Is it right??

OpenStudy (jhannybean):

No. this is 3 s.f. x 3 s.f = 3 s.f

OpenStudy (jhannybean):

\[\sf 1.82~ \frac{g}{mL} ~\times ~ 0.613~cm = 1.11566 ~\frac{g}{mL}\cdot cm = 1.12~\frac{g}{mL} \cdot cm\]

OpenStudy (mtalhahassan2):

wait how you get 0.613 cm

OpenStudy (jhannybean):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MTALHAHASSAN2 1.82 g/mL X 0.613 cm = 1.6X10^1 \(\color{#0cbb34}{\text{End of Quote}}\) I was basing it off what you wrote

OpenStudy (mtalhahassan2):

oh ok the question was thou 1.82 g/mL X 8.6 mL = 1.6 X 10 ^1

OpenStudy (mtalhahassan2):

sorry about that!!

OpenStudy (jhannybean):

okay, then your answer would be 2 s.f

OpenStudy (jhannybean):

\[\sf 1.82~\frac{g}{mL}~\times~8.6~mL~ =~ 15.652~g~ = ~16.~g\]